### v1

# Einstein manifolds of negative lower bounds on curvature operator of the second Kind

Abstract.
We demonstrate that $n$-dimension closed Einstein manifolds, whose smallest eigenvalue of the curvature operator of the second kind of $\mathring{R}$ satisfies $\lambda_{1}\geq-\theta(n)\bar{\lambda}$, are either flat or round spheres, where $\bar{\lambda}$ is the average of the eigenvalues of $\mathring{R}$, and $\theta(n)$ is defined as in equation (1.2). Our result improves a celebrated result (Theorem 1.1) concerning Einstein manifolds with nonnegative curvature operator of the second kind.
Key words and phrases:
Curvature operator of the second kind, Einstein manifolds, Sphere theorems
2020 Mathematics Subject Classification:
53C20, 53C24, 53C25
1. Introduction and main results
In this paper, we explore the sphere theorems for Einstein manifolds within the context of conditions imposed by the curvature operator of the second kind. The study of the curvature operator of the second kind, as defined in Section 2.2 below, can be traced back to a conjecture by Nishikawa [Nis86], which asserts that a closed Riemannian manifold with positive (or nonnegative) curvature operator of the second kind is diffeomorphic to a spherical space form (or a Riemannian locally symmetric space). Recently, Cao-Gursky-Tran [CGT23] and Li [Li24a] have confirmed Nishikawa’s conjecture for the positive and nonnegative cases, respectively. For further information and recent developments on this topic, we direct readers to the following references:[Li23a, Li22b, Li23b, Li24b, FL24, NPW23, NPWW23, DF24, DFY24].
For Einstein manifolds, a well-known result in this aspect is the following theorem.
Theorem 1.1.
A closed Einstein manifold with nonnegative curvature operator of the second kind is either flat or a round sphere.
[…]
Cao-Gursky-Tran [CGT23], Nienhaus-Petersen-Wink [NPW23], Li [Li24a] and Dai-Fu [DF24] have proved many excellent results under various curvature conditions. Recently, Li [Li24b] introduced new lower bound conditions on the curvature operator of the second kind (see [Li24b, Definition 1.2]) and proved optimal sphere theorems. Based on this framework, we focus on Li’s condition with $\alpha=1$, namely (1.1), which is weaker than nonnegative curvature operator of the second kind. To state our result, we introduce some notations: $(M^{n},g)$ is an $n$ dimensional Riemannian manifold, $\mathring{R}$ represents the curvature operator of the second kind, $N=(n+2)(n-1)/2$ is the dimension of $S_{0}^{2}(T_{p}M)$, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$. The main result of this paper is as follows:
Theorem 1.2.
Let $(M^{n},g)$ be a closed Einstein manifold of dimension $n\geq 4$. If
(1.1) $\lambda_{1}\geq-\theta(n)\bar{\lambda}.$
then $M$ is either flat or a round sphere. Here $\theta(n)$ is a positive constant given by
(1.2) $\begin{split}\theta(n)=\frac{3(N-1)(N+1-n)}{(N-1)(N-3)+3n(N-2)}-1.\end{split}$
Remark 1.1.
$\theta(n)$ defined by (1.2) is strictly positive and increasing in $n\in[4,\infty)$ and tends to 2 as $n\rightarrow\infty$.
Remark 1.2.
Theorem 1.2 essentially represents a gap theorem for Einstein manifolds, in the sense that if the curvature operator of the second kind satisfies (1.1), then the curvature operator of the second kind is nonnegative.
The crucial step in proving Theorem 1.2 is solving an extremum problem introduced in Proposition 3.4. Upon examination of the proof, it becomes evident that for any $\theta$, strictly greater than $\theta(n)$, Proposition 3.4 does not hold. Consequently, by employing our method, $\theta(n)$ as provided in equation (1.2) is deemed optimal. But it remains an interesting question that whether $\theta(n)$ is sharp or not for Theorem 1.2.
The subsequent sections of this paper are structured as follows. In Section 2, we give the definition of the curvature operator of the second kind and derive some identities for the curvatures on Einstein manifolds. Section 3 is devoted to the proof of Theorem 1.2.
2. Preliminaries
2.1. Notation and Conventions
In this section, we set up some notations and recall some basic facts about the curvature operator of the second kind, and refer to [CV60, BK78, OT79, Nis86, Kas93, Li24a] for more details.
Let $(M^{n},g)$ be an $n$-dimensional manifold, $V=T_{p}M$, $\{e_{i}\}_{i=1}^{n}$ be an orthonormal basis of $V$ with respect to $g$, and $S^{2}(V)$ be the space of symmetric two tensors on $V$. In this paper, we always identify $V$ with its dual space $V^{*}$ via the metric $g$. Then $S^{2}(V)$ splits into $O(V)$-irreducible subspaces as
$S^{2}(V)=S^{2}_{0}(V)\oplus\mathbb{R}g,$
where $S^{2}_{0}(V)$ is the space of traceless part of $S^{2}(V)$ and $g=\sum\limits_{i=1}^{n}e_{i}\otimes e_{i}$. We denote by $N$ the dimension of $S^{2}_{0}(V)$, that is
$N=\dim(S^{2}_{0}(V))=\tfrac{(n-1)(n+2)}{2}.$
The space of symmetric two-tensors on $\wedge^{2}V$, denoted by $S^{2}(\wedge^{2}V)$, has an orthogonal decomposition
$S^{2}(\wedge^{2}V)=S^{2}_{B}(\wedge^{2}V)\oplus\wedge^{4}V,$
where $S^{2}_{B}(\wedge^{2}V)$ is the space of algebraic curvature operators on $V$, consisting of all tensors $R\in S^{2}(\wedge^{2}V)$ satisfying the first Bianchi identity. The tensor product $\otimes$ is defined by
$(e_{i}\otimes e_{j})(e_{k},e_{l})=\delta_{ik}\delta_{jl},$
where
$\delta_{ij}=\begin{cases}1,&i=j,\\ 0,&i\neq j.\end{cases}$
For $e_{i}$ and $e_{j}$ in $V$, the symmetric product $\odot$ and wedge product $\wedge$ are defined by
$e_{i}\odot e_{j}=e_{i}\otimes e_{j}+e_{j}\otimes e_{i},$
and
$e_{i}\wedge e_{j}=e_{i}\otimes e_{j}-e_{j}\otimes e_{i},$
respectively. Furthermore the inner product on $\wedge^{2}(V)$ is defined by
$\langle A,B\rangle=\frac{1}{2}\operatorname{tr}(A^{T}B),$
and the inner product on $S^{2}(V)$ is defined by
$\langle A,B\rangle=\operatorname{tr}(A^{T}B).$
Clearly, $\{e_{i}\wedge e_{j}\}_{1\leq i<j\leq n}$ forms an orthonormal basis of $\wedge^{2}(V)$ and $\{\dfrac{1}{\sqrt{2}}e_{i}\odot e_{j}\}_{1\leq i<j\leq n}\cup\{\dfrac{1}{2}e_{i}\odot e_{i}\}_{1\leq i\leq n}$ forms an orthonormal basis of $S^{2}(V)$.
2.2. Curvature Operator of the Second Kind
For an algebraic curvature tensor $R\in S^{2}_{B}(\wedge^{2}(T_{p}M))$, $R$ can induce two self-adjoint operators (see [Nis86]). The first one denoted by $\hat{R}:\wedge^{2}(T_{p}M)\to\wedge^{2}(T_{p}M)$ is defined as
$\hat{R}(\omega)_{ij}=\frac{1}{2}\sum_{k,l=1}^{n}R_{ijkl}\omega_{kl},$
which is called the curvature operator of the first kind. The other one, $\overline{R}:S^{2}(T_{p}M)\to S^{2}(T_{p}M)$, is defined by
$\overline{R}(\varphi)_{ij}=\sum_{k,l=1}^{n}R_{iklj}\varphi_{kl}.$
The curvature operator of the second kind, introduced by Nishikawa [Nis86], is symmetric bilinear form
$\mathring{R}:S^{2}_{0}(T_{p}M)\times S^{2}_{0}(T_{p}M)\to\mathbb{R}$
obtained by restricting $\overline{R}$ to $S^{2}_{0}(T_{p}M)$, the space of traceless symmetric two-tensors. It was pointed out in [NPW23] that the curvature operator of the second kind $\mathring{R}$ can also be interpreted as the self-adjoint operator
(2.1) $\mathring{R}=\pi\circ\overline{R}:S^{2}_{0}(T_{p}M)\to S^{2}_{0}(T_{p}M),$
where $\pi:S^{2}(T_{p}M)\to S^{2}_{0}(T_{p}M)$ is the projection map.
We denote by $\{\lambda_{j}\}_{j=1}^{N}$ the eigenvalues of $\mathring{R}$, by $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ be the average of the eigenvalues of $\mathring{R}$, by $R$ the Riemann curvature tensor, and by ${\operatorname{Scal}}$ be the scalar curvature. For Einstein manifolds, one can obtain the scalar curvature via $\mathring{R}$. Precisely
Proposition 2.1.
Suppose $(M^{n},g)$ is an Einstein manifold of dimension $n$. Then
(2.2) ${\operatorname{Scal}}=n(n-1)\bar{\lambda}.$
Proof.
Let $\{e_{i}\}_{i=1}^{n}$ be an orthonormal basis of $T_{p}M$, then $\{\dfrac{1}{\sqrt{2}}e_{i}\odot e_{j}\}_{1\leq i<j\leq n}\cup\{\dfrac{1}{2}e_{i}\odot e_{i}\}_{1\leq i\leq n}$ is an orthonormal basis of $S^{2}(T_{p}M)$. Since
$\left\langle{\bar{R}(\frac{1}{{\sqrt{2}}}{e_{i}}\odot{e_{j}}),\frac{1}{{\sqrt{2}}}{e_{i}}\odot{e_{j}}}\right\rangle=R_{ijij},$
then
$\operatorname{tr}(\bar{R})=$ $\sum\limits_{1\leq i<j\leq n}{\left\langle{\bar{R}(\frac{1}{{\sqrt{2}}}{e_{i}}\odot{e_{j}}),\frac{1}{{\sqrt{2}}}{e_{i}}\odot{e_{j}}}\right\rangle}+\sum\limits_{1\leq i\leq n}{\left\langle{\bar{R}(\frac{1}{2}{e_{i}}\odot{e_{i}}),\frac{1}{2}{e_{i}}\odot{e_{i}}}\right\rangle}$
$=$ $\sum\limits_{1\leq i<j\leq n}R_{ijij}$
$=$ $\frac{{\operatorname{Scal}}}{2}.$
Thus,
$\operatorname{tr}(\mathring{R})=\operatorname{tr}(\bar{R})-\langle\bar{R}(\frac{g}{\sqrt{n}}),\frac{g}{\sqrt{n}}\rangle=\frac{{\operatorname{Scal}}}{2}+\frac{{\operatorname{Scal}}}{n}=\frac{n+2}{2n}{\operatorname{Scal}}.$
Using $\operatorname{tr}(\mathring{R})=N\bar{\lambda}$, we obtain
${\operatorname{Scal}}=\frac{2n}{n+2}N\bar{\lambda}=n(n-1)\bar{\lambda},$
proving (2.2). $\square$
Definition 2.2 ([NPW23]).
Let $\mathcal{T}^{(0,k)}(V)$ denote the space of $(0,k)$-tensor space on $V$. For $S\in S^{2}(V)$ and $T\in\mathcal{T}^{(0,k)}(V)$, we define
$S:$ $\mathcal{T}^{(0,k)}(V)\to\mathcal{T}^{(0,k)}(V),$
$(ST)({X_{1}},\cdots,{X_{k}})=\sum\limits_{i=1}^{k}{T({X_{1}},\cdots,S{X_{i}},\cdots,{X_{k}})},$
and define $T^{S^{2}}\in\mathcal{T}^{(0,k)}(V)\otimes S^{2}(V)$ by
$\left\langle T^{S^{2}}(X_{1},\cdots,X_{k}),S\right\rangle=(ST)\left(X_{1},\cdots,X_{k}\right).$
According to the above definition, if $\{\bar{S}^{j}\}_{j=1}^{N}$ is an orthonormal basis for $S^{2}(V)$, then
$T^{S^{2}}=\sum\limits_{j=1}^{N}\bar{S}^{j}T\otimes\bar{S}^{j}.$
Similarly, we define $T^{S^{2}_{0}}\in\mathcal{T}^{(0,k)}(V)\otimes S^{2}_{0}(V)$ by
$T^{S^{2}_{0}}=\sum\limits_{j=1}^{N}S^{j}T\otimes S^{j},$
where $\{S^{j}\}_{j=1}^{N}$ is an orthonormal basis for $S^{2}_{0}(V)$.
2.3. A formula for Weyl tensor on Einstein manifolds
Given $A,B\in S^{2}(V)$, their Kulkarni-Nomizu product gives rise to $A\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}B\in S^{2}_{B}(\wedge^{2}V)$ via
$(A\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}B)_{ijkl}=A_{ik}B_{jl}+A_{jl}B_{ik}-A_{jk}B_{il}-A_{il}B_{jk}.$
It is well known that the Riemann curvature tensor $R$ can be decomposed into irreducible components (cf. [CM20, (1.79)]) as
(2.19) $R=W+\frac{1}{n-2}\operatorname{Ric}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g-\frac{{\operatorname{Scal}}}{2(n-1)(n-2)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g,$
where $W$ is the Weyl curvature tensor and $\operatorname{Ric}$ is the Ricci curvature tensor. In terms of any basis, we have
$R_{ijkl}=$ $W_{ijkl}+\frac{1}{n-2}\left(R_{ik}g_{jl}+R_{jl}g_{ik}-R_{il}g_{jk}-R_{jk}g_{il}\right)$
$-\frac{{\operatorname{Scal}}}{(n-1)(n-2)}\left(g_{ik}g_{jl}-g_{il}g_{jk}\right).$
Moreover, if $(M^{n},g)$ is an Einstein manifold, then $\operatorname{Ric}=\frac{{\operatorname{Scal}}}{n}g$, and equation (2.19) simplifies to
(2.28) $R=W+\frac{{\operatorname{Scal}}}{2n(n-2)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g.$
3. Proof of Theorem 1.2
We recall some curvature notations on $(M^{n},g)$: $R$ represents the Riemann curvature operator, $W$ represents the Weyl tensor, $\mathring{R}$ represents the curvature operator of the second kind, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$ To prove the main theorem, we conduct several computations involving the Riemann curvature through the curvature operator of the second kind on Einstein manifolds.
Lemma 3.1.
Let $(M^{n},g)$ be an $n$ dimensional Einstein manifold. Then
(3.1) $|R|^{2}=|W|^{2}+2n(n-1)\bar{\lambda}^{2},$
and
(3.2) $\sum_{j=1}^{N}\lambda_{j}^{2}=\frac{3}{4}\left|R\right|^{2}-(n-1)^{2}\bar{\lambda}^{2}.$
Proof.
Recall from the equation (2.28) that
$R=\frac{{\operatorname{Scal}}}{2n(n-1)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g+W,$
then
$|R|^{2}=\frac{{\operatorname{Scal}}^{2}}{4n^{2}(n-1)^{2}}|g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g|^{2}+|W|^{2}=\frac{{\operatorname{Scal}}^{2}}{2n(n-1)}+|W|^{2},$
where we used the identity $|g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g|^{2}=8n(n-1)$ (cf. [Lee18, Lemma 7.22]). Combining this with identity (2.2), we conclude
$|R|^{2}=|W|^{2}+2n(n-1)\bar{\lambda}^{2},$
giving (3.1).
Recall from Lemma 3.3 of [DF24] that
$\sum_{j=1}^{N}\lambda_{j}^{2}=\frac{3}{4}\left|R\right|^{2}-\frac{{\operatorname{Scal}}^{2}}{n^{2}},$
and use identity (2.2) again, we have
$\sum_{j=1}^{N}\lambda_{j}^{2}=\frac{3}{4}\left|R\right|^{2}-(n-1)^{2}\bar{\lambda}^{2},$
giving (3.2). $\square$
Lemma 3.2.
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Suppose the smallest eigenvalue of $\mathring{R}$ satisfies $\lambda_{1}\geq-\theta\bar{\lambda}$, then
(3.3) $\sum_{j=1}^{N}\lambda_{j}|S^{j}W|^{2}\geq-\frac{8(n^{2}+n-8)}{3n}\theta\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+\frac{4(n^{2}+n-8)(n-1)(n+2)}{3n}\theta\bar{\lambda}^{3},$
where $S^{j}W$ is given by Definition 2.2.
Proof.
Recall from equality (4.1) of [DF24] that
$\sum_{j=1}^{N}|S^{j}W|^{2}=\frac{2(n^{2}+n-8)}{n}|W|^{2},$
and by the assumption $\lambda_{1}\geq-\theta\bar{\lambda}$, we estimate that
(3.4) $\sum_{j=1}^{N}\lambda_{j}|S^{j}W|^{2}\geq-\theta\bar{\lambda}\sum_{j=1}^{N}|S^{j}W|^{2}=-\theta\bar{\lambda}\frac{2(n^{2}+n-8)}{n}|W|^{2}.$
From (3.1) and (3.2), we have
$|W|^{2}=\frac{4}{3}\sum_{j=1}^{N}\lambda_{j}^{2}-\frac{2}{3}(n-1)(n+2)\bar{\lambda}^{2},$
and substituting above equality into (3.4), we conclude inequality (3.3) holds. $\square$
Lemma 3.3.
Suppose $(M^{n},g)$ is an Einstein manifold satisfying $\lambda_{1}\geq-\theta\bar{\lambda}$. Then
(3.5) $\begin{split}\frac{9n}{16}\langle\Delta R,R\rangle\geq&\Big{[}N(N-3)\theta-(2N-9n+6)N\Big{]}\bar{\lambda}^{3}\\ &+\Big{[}(2N-12n+6)-(N-3)\theta\Big{]}\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+3n\sum_{j=1}^{N}\lambda_{j}^{3},\end{split}$
where $\Delta$ is the Laplace-Beltrami operator with respect to the metric $g$.
Proof.
Recall from equality (3.3) of [DF24] that
$3\langle\Delta R,R\rangle=$ $\displaystyle\sum_{j=1}^{N}{{\lambda_{j}}{{\left|{{S^{j}}W}\right|}^{2}}}+8\left(\frac{-n^{3}+6n^{2}+12n-8}{3n^{4}(n-1)^{2}}\right){{\operatorname{Scal}}^{3}}$
$+8\left(\frac{2n^{2}-22n+8}{3n^{2}(n-1)}\right){\operatorname{Scal}}\displaystyle\sum_{j=1}^{N}{\lambda_{j}^{2}}+16\displaystyle\sum_{j=1}^{N}{\lambda_{j}^{3}},$
and using ${\operatorname{Scal}}=n(n-1)\bar{\lambda}$, we obtain
$3\langle\Delta R,R\rangle=$ $\sum_{j=1}^{N}{{\lambda_{j}}{{\left|{{S^{j}}W}\right|}^{2}}}+\frac{8(n-1)}{3n}\left(-n^{3}+6n^{2}+12n-8\right){\bar{\lambda}^{3}}$
$+8\left(\frac{2n^{2}-22n+8}{3n}\right)\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+16\sum_{j=1}^{N}{\lambda_{j}^{3}}.$
By (3.3) and $N=(n-1)(n+2)/2$, we derive that
$3\langle\Delta R,R\rangle\geq$ $-\frac{8(n^{2}+n-8)\theta}{3n}\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+\frac{4(n^{2}+n-8)(n-1)(n+2)}{3n}\theta\bar{\lambda}^{3}+16\sum_{j=1}^{N}{\lambda_{j}^{3}}$
$+\frac{8(n-1)}{3n}\left(-n^{3}+6n^{2}+12n-8\right){\bar{\lambda}^{3}}+8\left(\frac{2n^{2}-22n+8}{3n}\right)\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}$
=$\quad$= $\frac{16}{3n}\Big{[}N(N-3)\theta-(2N-9n+6)N\Big{]}\bar{\lambda}^{3}$
$+\frac{16}{3n}\Big{[}(2N-12n+6)-(N-3)\theta\Big{]}\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+16\sum_{j=1}^{N}\lambda_{j}^{3},$
proving (3.5). $\square$
Let $\lambda:=(\lambda_{1},\lambda_{2},\cdots,\lambda_{N})$, and we denote the right hand side of (3.5) by $f(\lambda)$, namely
$\begin{split}f(\lambda):=&\Big{[}N(N-3)\theta-(2N-9n+6)N\Big{]}\bar{\lambda}^{3}\\ &+\Big{[}(2N-12n+6)-(N-3)\theta\Big{]}\bar{\lambda}\sum_{j=1}^{N}\lambda_{j}^{2}+3n\sum_{j=1}^{N}\lambda_{j}^{3}.\end{split}$
Proposition 3.4.
Let $\theta$ be the positive constant defined by (1.2). If $\bar{\lambda}>0$ and $-\theta\bar{\lambda}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$, then
$f(\lambda)\geq 0,$
and the equality holds if and only if
$\lambda=(1,1,\cdots,1)\bar{\lambda}\text{\quad or \quad}\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}.$
Proof.
Let $\tilde{\lambda}_{j}=\lambda_{j}/\bar{\lambda}$, $\tilde{\lambda}=(\tilde{\lambda}_{1},\tilde{\lambda}_{2},\cdots,\tilde{\lambda}_{N})$, and then $\tilde{\lambda}$ satisfies the constraints
(3.6) $\sum_{j=1}^{N}\tilde{\lambda}_{j}=N,\quad\text{ and }\quad\tilde{\lambda}_{j}\geq-\theta.$
Set $x=(\tilde{\lambda}_{1}+\theta,\tilde{\lambda}_{2}+\theta,\cdots,\tilde{\lambda}_{N}+\theta)$ and $t=1+\theta$, then the constraints (3.6) are equivalent to
$\sum_{j=1}^{N}x_{j}=Nt\quad\text{and}\quad x_{j}\geq 0.$
From the expression of $f(\lambda)$ we can deduce that
$\frac{1}{\bar{\lambda}^{3}}f(\lambda)=C(n,t)+3nF(x),$
where
$C(n,t)=$ $N\big{[}(N-3)t-(3N-9n+3)\big{]}+3nN(t-1)^{2}(2t+1)$
$+\big{[}3(N-4n+1)-(N-3)t\big{]}\big{(}-2Nt(t-1)+N(t-1)^{2}\big{)},$
and
$F(x)=-(3-\frac{N-2}{N-1})t\sum_{j=1}^{N}x_{j}^{2}+\sum_{j=1}^{N}x_{j}^{3}.$
Now we reduce the proposition to seek the minimal points of $F(x)$ with constraints
$\sum_{j=1}^{N}x_{j}=Nt,\quad\text{and}\quad x_{j}\geq 0\ (j=1,2,\cdots N).$
The corresponding Lagrangian function is
$\mathcal{L}(x,\mu):=-(3-\frac{N-2}{N-1})t\sum_{j=1}^{N}x_{j}^{2}+\sum_{j=1}^{N}x_{j}^{3}+\mu(\sum_{j=1}^{N}x_{j}-Nt).$
Let
$\Omega_{N}=\Big{\{}x\in\mathbb{R}^{N}:\sum_{j=1}^{N}x_{j}=Nt,x_{j}>0\Big{\}},$
and
$\Omega_{k}=\Big{\{}x\in\mathbb{R}^{N}:\sum_{j=1}^{N}x_{j}=Nt,x_{1}=\cdots=x_{N-k}=0,x_{j}>0,j\geq N-k+1\Big{\}}$
for $k=1,\cdots,N-1$.
In $\Omega_{k}$, using the method of Lagrange multipliers, we find that the critical points
$x=\big{(}\underbrace{0,\cdots,0}_{N-k},x_{N-k+1},\cdots,x_{N}\big{)}$
of $\mathcal{L}(x,\mu)$ satisfy
$\begin{cases}3x_{i}^{2}-2(3-\frac{N-2}{N-1})tx_{i}+\mu=0,\\ \sum_{i=N-k+1}^{N}x_{i}=Nt,\end{cases}$
hence the possible critical points are given by
$P_{k,l}:=\big{(}{\underbrace{0,\cdots,0}_{N-k},\underbrace{a,\cdots,a}_{l},\underbrace{b,\cdots,b}_{k-l}}\big{)},$
for $0\leq l\leq\frac{k}{2}$, where $a$ and $b$ are defined through
(3.7) $\begin{cases}a+b=2(1-\frac{N-2}{3(N-1)})t,\\ la+(k-l)b=Nt.\end{cases}$
When $l$ equals $\frac{k}{2}$ and equation (3.7) is solvable, we have $a+b=\frac{2N}{k}t$ and $1-\frac{N-2}{3(N-1)}=\frac{N}{k}$, and in this case direct calculation reveals that
(3.8) $F\big{(}{\underbrace{0,\cdots,0}_{N-k},\underbrace{a,\cdots,a}_{k/2},\underbrace{b,\cdots,b}_{k/2}}\big{)}=F(P_{k,0})$
for $\forall a,b\geq 0$ satisfying $a+b=2Nt/k$.
When $l<\frac{k}{2}$, by solving (3.7) we derive
$a=(A-B)t,\quad b=(A+B)t,$
where $A=1-\frac{N-2}{3(N-1)}$, $B=\frac{N-Ak}{k-2l}$, and direct calculation gives
$F(P_{k,l})=t^{3}\left(-2kA^{3}-3A^{2}(N-Ak)+\frac{(N-Ak)^{3}}{(k-2l)^{2}}\right),$
hence
(3.9) $F(P_{k,0})<F(P_{k,l}),\quad 1\leq l<k/2.$
Therefore we conclude from (3.8) and (3.9) that
$\min\Big{\{}F(x):\sum_{j=1}^{n}x_{j}=Nt,x_{j}\geq 0\Big{\}}=\min\Big{\{}F(P_{k,0}),k=1,2,\cdots,N\Big{\}},$
which indicates that the minimum points of the function $f(\lambda)$, subject to the constraint $\lambda_{i}\geq-\theta\bar{\lambda}$, are included among
$\lambda^{m}=(\underbrace{-\theta\bar{\lambda},\cdots,-\theta\bar{\lambda}}_{m},\underbrace{\frac{N+m\theta}{N-m}\bar{\lambda},\cdots,\frac{N+m\theta}{N-m}\bar{\lambda}}_{N-m})$
for $m=0,1,\cdots,N-1$.
For $\lambda=\lambda^{m}$, we calculate that
$\frac{1}{\bar{\lambda}^{2}}\sum_{j=1}^{N}\lambda^{2}_{j}=m\theta^{2}+(N-m)\left(\frac{N+m\theta}{N-m}\right)^{2}=N+\frac{N}{N-m}mt^{2},$
and
$\frac{1}{\bar{\lambda}^{3}}\sum_{j=1}^{N}\lambda^{3}_{j}=-mt^{3}\frac{(N-2m)N}{(N-m)^{2}}+\frac{3mN}{N-m}t^{2}+N,$
we then conclude
(3.10) $\frac{1}{\bar{\lambda}^{3}}f(\lambda^{m})=$ $\Big{[}-\frac{N(N-3)m}{N-m}-\frac{3nmN(N-2m)}{(N-m)^{2}}\Big{]}t^{3}+\frac{3N(N+1-n)}{N-m}mt^{2}$
$=$ $\frac{mNt^{2}}{N-m}\Big{[}3(N+1-n)-(N-3+\frac{3n(N-2m)}{N-m})t\Big{]}.$
From the definition of $\theta$ in (1.2), we see
$t=1+\theta=\frac{3(N+1-n)}{N-3+3n(N-2)/(N-1)},$
and substituting above to (3.10), we find
$f(\lambda^{0})=f(\lambda^{1})=0,\quad\text{and}\quad f(\lambda^{k})>f(\lambda^{1})=0\quad(k=2,3,\cdots,N-1).$
Hence Proposition 3.4 comes true. $\square$
Proof of the Theorem 1.2.
If $\bar{\lambda}>0$, then Lemma 3.3 and Proposition 3.4 imply
(3.11) $\langle\Delta R,R\rangle\geq 0.$
If $\bar{\lambda}=0$, $\lambda_{1}\geq-\theta(n)\bar{\lambda}$ implies $\lambda_{1}=\lambda_{2}=\cdots\lambda_{N}=0$, and (3.11) holds by Lemma 3.3. Using (3.11), we estimate that
$\Delta|R|^{2}=2|\nabla R|^{2}+2\langle\Delta R,R\rangle\geq 2|\nabla R|^{2},$
and integrating above inequality over $M$ yields
[…]
which implies $|\nabla R|=0$. Therefore $M$ is a symmetric space. [Li24b, Proposition 2.8] and [Li24b, Theorem 1.8(2)] imply that $M$ is either flat or a rational homology sphere. Besides, Poincaré duality shows that rational cohomology is determined by rational homology [Hat02, Chapter 3.3]. Thus we conclude that $M$ is either flat or a symmetric space which is also a rational cohomology sphere. Compact symmetric spaces which are real cohomology spheres were classified by Wolf in [Wol69, Theorem 1]. If $\bar{\lambda}>0$, apart from spheres, $SU(3)/SO(3)$ is the only simply connected example. Proposition 3.4 implies that the eigenvalues of $\mathring{R}$ is either $\lambda=(1,\cdots,1)\bar{\lambda}$ or $\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}$, and combining with [NPW23, Example 4.5], we know that $M$ is a round sphere. $\square$
Acknowledgments
The first author expresses her gratitude to her advisor, Professor Ying Zhang, for lots of encouragement and helpful suggestions. The authors would like to thank Professor Xiaolong Li for many helpful discussions. The research of this paper is partially supported by NSF Grant No. 12171345 and NSF of Jiangsu Province Grant No. BK20231309.